OpenStudy (anonymous):

help me create an example? i did the word part but i need an example for each

OpenStudy (anonymous):

here are the questions with my answers: 1. If I know the lengths of two sides of a right triangle, how do I find the third? Pythagorean Theorem; a^2 + b^2 = c^2 2. Could I find the two missing side lengths of a right triangle if I only know one side length and one angle measure (other than the 90 degree angle)? yes. Law of Sines; sin(a)/A = sin(b)/B = sin(c)/C where a,b,c are the angles of the triangle and A, B, C are the projected sides. 3. Could I find the two missing angle measures if I know some of the side lengths of a right triangle? Use the theorem to find the lengths of triangle and apply them to the Law of Sines. You will need to have at least 2 lengths known. 4. What makes a triangle a “special” right triangle? How can special right triangles help me find side lengths? Special right triangles have set ratios for the lengths of each side of the triangle.

OpenStudy (phi):

In 4. are they asking about 30-60-90 or 45-45-90 triangles? Examples for 1) use the standard 3-4-5 right triangle. label two of the legs and mark the third unknown

OpenStudy (anonymous):

in number four it doesnt spicifically state anything

OpenStudy (phi):

You can use a 30-60-90 as an example in 4. label the longest side 1, the side opposite 30 = 1/2 and the remaining side sqrt(3)/2

OpenStudy (phi):

For 2. use a 30-60-90 triangle. label the 30 deg angle as known, and the side opposite it as 1/2 show you can find the hypotenuse: h/sin(90) = 0.5 / sin(30) and show you can find the 2nd side: x/sin(60) = 0.5/ sin(30)

OpenStudy (phi):

for 3) label the two legs 3 and 4. (that will make the hyp 5) then use sin(A)/3 = sin(90)/5 to find the A opposite the side with length 3 sin(B)/4 = sin(90)/5 to find the angle B opposite the side with length 4

OpenStudy (anonymous):

so for number one just put this,

OpenStudy (anonymous):



OpenStudy (phi):

yes. then label h . then write 3^2+4^2= h^2 9+16= h^2 , 25=h^2 , 5= h

OpenStudy (anonymous):

so this will be my drawing and then just right the equation you wrote